# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v1

@@ 4 Main Method
In this section, we introduce circuit-tuning, an algorithm based on the theory of node redundancy for both fine-tuning neural networks and analyzing training dynamics.
@@
  Input: dataset $\mathcal{X}$, model $M$, loss function $\mathcal{L}$, the metric for measuring indirect effect $\mathcal{L}_{m}$, the number of edges to save $N$, the optimization steps after circuit discovery $K$.
  Initialize the computational graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ from $M$, the circuit $\mathcal{C}=\mathcal{G}$, and the iteration step $i=0$.
@@
           Intervene: $n_{2}(x^{\prime})\!=\!n_{2}^{direct}\big{(}x|do\big{(}n_{1}\!\leftarrow n_{1}\!(x^{\prime})\big{)}\big{)}$
           Edge contribution: $c(e)=\mathbb{E}_{x_{i}\sim\mathcal{X}_{T}}\big{[}|IE(e;x_{i})|\big{]}$
@@
        $\mathcal{V}_{T}\leftarrow\!\{n\,|\,n\!\in\!\mathcal{V}\land n\text{ is incident to an edge }e\!\in\!\mathcal{E}_{T}\}$
        $\mathcal{C}\leftarrow(\mathcal{V}_{T},\mathcal{E}_{T})$
@@
     Update the parameters in $\mathcal{C}$
@@
4.1 Circuit-tuning
4.1.1 From Nodes to Edges
The definition of node redundancy as well as the detection of it is provided in Section 3.3. Since the nodes inside a graph are connected by edges, we can also study the node redundancy from the view of edges.
Following the definition of the attribution score in edge attribution patching (EAP) (Syed et al., 2023) and the discussions in Section 3.3, we can define the contribution of an edge $e$ to the output as follows:
$c(e)$ $=\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}|c_{i}(e)|\big{]}$ (5)
$=\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}|IE(e;x_{i})|\big{]}$
where the edge $e:n_{1}\rightarrow n_{2}$ is connected with an upstream node $n_{1}$ and a downstream node $n_{2}$. To measure the direct effect from $n_{1}$ to $n_{2}$, we set the value of $n_{1}$ to another value $n_{1}(x^{\prime})$ while keeping all the other nodes between $n_{1}$ and $n_{2}$ unchanged. Then the affected value of $n_{2}$ is:
$n_{2}(x^{\prime})=n_{2}(x)-n_{2}^{n_{1}}(x)+n_{2}^{n_{1}}(x^{\prime})$
where $n_{2}^{n_{1}}(\cdot)$ denotes the contribution from $n_{1}$ to $n_{2}$, and
$n_{2}^{n_{1}}(x^{\prime})=n_{2}^{n_{1}}\big{(}x\,|\,do\big{(}n_{1}\leftarrow n_{1}(x^{\prime})\big{)}\big{)}$
Then the indirect effect from $e$ to the final output is
$\begin{split}IE(e;x)=&IE(n_{1}\rightarrow n_{2};x)\\ =&\mathcal{L}_{m}\big{[}M\big{(}x\,|\,do\big{(}n_{2}\leftarrow n_{2}(x^{{\prime}})\big{)}\big{)}\big{]}-\mathcal{L}_{m}\big{[}M(x)\big{]}\end{split}$
To reduce computation, we apply a first-order Taylor expansion to $n_{1}^{\prime}$ and $n_{2}^{\prime}$ respectively. Then the calculation of edge contribution requires only one forward and one backward pass. See Appendix A for derivation details.
Therefore the approximation of contribution $c(e)$ is available which can be used to measure the importance of an edge. We can determine the redundancy of a node by checking the contributions of the edges connected to it.
4.1.2 The Circuit-tuning Algorithm
Given a model $M$ for fine-tuning, the model is initialized into a computational graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ in which $\mathcal{V}$ and $\mathcal{E}$ represent the nodes and edges in $\mathcal{G}$ respectively. A metric $\mathcal{L}_{m}$ and a loss function $\mathcal{L}$ are used for circuit discovery and parameter optimization respectively. Given a dataset $\mathcal{X}\sim\mathcal{D}_{T}$ for the fine-tuning task $T$, circuit-tuning alternately performs the following two procedures:
Circuit Discovery For a batch of data $\mathcal{X}_{T}\in\mathcal{X}$, we calculate the contribution of each edge $e\in\mathcal{E}$ according to (5). Then all the edges are sorted in descending order based on their contributions, and the edges with top $N$ contributions are selected. Then the graph $\mathcal{G}$ is pruned into a circuit $\mathcal{C}=(\mathcal{V}_{T},\mathcal{E}_{T})$ with only the selected edges $\mathcal{E}_{T}$ together with the nodes $\mathcal{V}_{T}$ at both ends of each edge inside.
Circuit-tuning For a batch of data $\mathcal{X}_{T}\in\mathcal{X}$, all the parameters outside $\mathcal{C}$ are frozen, and only the parameters corresponding to the nodes $\mathcal{V}_{T}$ inside $\mathcal{C}$ are updated through a forward pass and a backward pass as usual. After $K$ steps of optimization, all the frozen parameters are freed and the graph $\mathcal{G}$ is reset to its original state.
The full process is shown in Algorithm 1. The parameter to optimize corresponds to the parameter matrix that maps an input to the activation of a node in the circuit. For example, the parameter corresponding to the node at the attention output refers to $W_{O}$, which is the output projection matrix.
@@ 4.2 Characteristics of Our Method
Capability Acquisition Different from recent studies that focus on the circuits of certain model behaviors, circuit-tuning aims to automatically find the relevant components responsible for a specific task and guide the model to develop an ability through parameter optimization. The dynamics of this process will be verified in Section 5.1.
@@
(c) logit difference after revision
Figure 1: Results in the subject-verb disagreement task
Precise Fine-tuning Compared with full fine-tuning and LoRA-based methods, our method only update a few parameters instead of taking all of the parameters into account. Since our method tries to leave out the irrelevant structures in terms of a target task, the general abilities are expected not to be affected, which will be verified in Section 5.
4.3 Convergence Analysis
To demonstrate the stability of circuit-tuning, we provide an analysis of the training convergence. Specifically, let’s consider the model as a continuously differentiable function $f(\theta)$. Let $\theta\in\mathbb{R}^{D}=(\theta_{1},\theta_{2},...,\theta_{D})$, and suppose that $f$ satisfies the conditions for Lipschitz gradient continuity:
$||\nabla_{\theta}f(\theta^{\prime})-\nabla_{\theta}f(\theta)||\leq L||\theta^{\prime}-\theta||$
where $L$ is a Lipschitz constant.
If we mask out a set of parameters $\theta_{mask}\subset\theta$, then we have
$\begin{split}&||\nabla_{\theta\backslash\theta_{mask}}f(\theta^{\prime})-\nabla_{\theta\backslash\theta_{mask}}f(\theta)||\\ <&||\nabla_{\theta}f(\theta^{\prime})-\nabla_{\theta}f(\theta)||\\ \leq&L||\theta^{\prime}-\theta||\end{split}$
This is because for each masked parameter $\theta_{i}\in\theta_{mask}$, the gradient of it equals to 0, which leads to a smaller norm for gradient change. Thus a smaller Lipschitz constant $L_{mask}<L$ serves as the upper bound for the change rate of gradients, which means the training process will be more stable compared with full fine-tuning. The evidence for this conclusion will be discussed in Section 5.1.3.
5 Experiments
In this section, we test our method across a variety of models and tasks. To comprehensively investigate the performance as well as the training dynamics during fine-tuning, we arrange the experiments into two parts. The first part serves as the verification and analysis of our method, while the second part aims to show that our method is scalable to various model sizes and tasks.
5.1 The Subject-verb Disagreement Task
5.1.1 Task Description and Data Preparation
The goal of the subject-verb disagreement task is to match a verb with a subject in an abnormal way. For example, “I is”, “he are” and “the cows eats” are all expected results for this task. In each sentence, the word before the verb is called the END token. The automatic evaluation metric for this task is the logit difference between the flipped verb and the original verb at the END token. This is because the logit at the END token is directly used for predicting the verb.
The reason why we create this task as the start of our experiments is that we expect the model to acquire a new capability from scratch. Besides, the task is simple and interesting for analyzing the training dynamics during fine-tuning.
For training data, we use the first 10k samples in the Pile (Gao et al., 2020) corpus. We extract 30k sentences in the present tense in English and flip the forms of the verbs in each sentence. For details, please refer to Section C.1.
5.1.2 Implementation Details
We use GPT2-small (Radford et al., 2019) in this task. We set the output of the attention and the MLP in each layer as upstream nodes, and the input of the query, key, value and the MLP in each layer as downstream nodes. During training, we use the logit difference discussed before as the metric $\mathcal{L}_{m}$ and follow (Syed et al., 2023) to calculate the edge contributions. We sweep over a range of $N$ which is the number of edges to be saved during circuit discovery. Mini-batch SGD is used for optimization. For details of settings, please refer to Section C.2.
@@ 5.1.3 Main Results
From Figure 1(a), we observe a flip in logit difference from negative to positive, which means the model adjusts its grammar to fit the data distribution of subject-verb disagreement. One noteworthy finding is that the trained model can generate abnormal texts in the past tense, such as “to be or not to be, that were a …”, which implies that the model really learns the new grammar and applies it smartly, since there is no sentence in the present tense in our training data at all.
@@
We find that with the increasing number of top $N$ edges, the logit difference gets bigger until $N$ reaches around 1000. We check and find that the number of tunable parameters for each $N$ is also saturated at this point. We believe when $N=1000$, almost all the necessary parameters for this task are included, and the performance cannot be further improved, suggesting the existence of intrinsic dimension. For verification of this, please refer to Section C.3.
@@
In Figure 1(b), we find that the perplexity (PPL) of full fine-tuning is high and fluctuates wildly during training, though the logit difference of it is higher. This observation implies better training stability as well as better preservation of general capabilities of circuit-tuning over full fine-tuning.
5.1.4 The Quality of the Discovered Circuit
To demonstrate that our method is able to find the required parameters accurately, we (i) calculate the faithfulness and completeness (Marks et al., 2024) of the circuits and (ii) provide an ablation study in which we randomly unfreeze the nodes outside the circuit to check their influence on the performance of this task. Results show from various aspects that the parameters required are found accurately. For details, please refer to Section C.3.
5.1.5 Analyses on Training Dynamics
We analyze the circuits before, during, and after training, and report some amazing findings that are worth studying.
(a) Before fine-tuning
(b) After fine-tuning
Figure 2: The flip in the Subject Attribute Heads. Each square corresponds to the $j$-th attention head in layer $i$, in which $i$ and $j$ correspond to the vertical and horizontal axis respectively.
@@ 5.1.5 Analyses on Training Dynamics
Interpretation of the Circuits To interpret the circuit into human-understandable structures, we follow (Wang et al., 2022) and decompose the circuit into heads of different functions. Firstly we find out the attention head that directly affects the output, which is called the Subject Attribute Head. Then we find out the head responsible for the localization of the subject, which is called the Subject Identification Head. Finally, we find out the Collaborative Heads that could affect the behaviors of the Subject Attribute Heads. For details of the processes above, please refer to Section C.4.1.
@@
The Flip of the Subject Attribute Heads The Subject Attribute Heads are responsible for matching the subject and the verb. We visualize the heads before and after fine-tuning in Figure 2. Through comparison, we observe an obvious flip at head.10.9, which implies the reversal in its function from subject-verb agreement to disagreement. Other heads (3.0, 6.0, 4.4, 11.8, etc) also see flips with varying degrees, which serves as strong evidence for the self-adjustment inside the nodes.
@@
The Sharing of the Subject Identification Heads The Subject Identification Heads attend heavily to the subject. One type of these heads attends to the END token (0.1, 0.3, etc), which is helpful to the cases like “they are”; the other type of heads (8.5, 10.5, etc) attends to the subject several tokens before, which is helpful to the cases like “the girl wearing a dress is”. Both types of heads perform the same before and after fine-tuning, which means their functions are preserved and shared all the way down.
@@
The Evidence of Hebbian Learning We visualize the circuits during the fine-tuning process in Figure 9, and observe a phenomenon akin to Hebbian learning (Do, 1949). We find that several edges are strengthened during training, just as the synapses between neurons can be strengthened after continuous stimulation. We regard this finding as evidence of the self-organization (Kohonen, 2012) inside the model to a new environment. See Section C.4.3 for details.
@@
Due to limited space, all the above together with other interesting findings are discussed in detail in Section C.4.
5.1.6 Improvement on Attribution Patching

### v2

@@ 4 Main method
4.1 Circuit-tuning
From nodes to edges Following the discussions in Section 3, we introduce circuit-tuning, an interpretable algorithm that allows transparent and precise fine-tuning. The only difference is that we shift our view from nodes to edges. The reason is that (i) we are interested in the interactions between nodes and (ii) the redundancy in nodes can be identified indirectly by checking the contributions of the edges connected to it. The shift from edges to nodes is actually an extension of the estimation method in Section 3.3, and the underlying idea is the same. Thus we can leverage the path patching technique in circuit discovery to identify the intrinsic dimensions. Discussions are detailed in Appendix D.
Given a model $M$ for fine-tuning, the model is initialized into a computational graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, as defined in Section 3.2. A metric $\mathcal{L}_{m}$ and a loss function $\mathcal{L}$ are used for output quantification and parameter optimization respectively, as described in Section 3.3. Given a dataset $\mathcal{X}\sim\mathcal{D}_{T}$ for the fine-tuning task $T$, circuit-tuning alternately performs the following two procedures:
(1) Graph pruning: For a batch of data $\mathcal{X}_{T}\in\mathcal{X}$, we compute the edge contribution for each edge $e:n_{1}\rightarrow n_{2}\in\mathcal{E}$ according to Equation (D), where $n_{1}$ and $n_{2}$ are the upstream and downstream nodes respectively. Then all the edges are sorted in descending order based on their contributions, and the edges with top $N$ contributions are selected. Then the graph $\mathcal{G}$ is pruned into a circuit $\mathcal{C}=\{\mathcal{V}_{T},\mathcal{E}_{T}\}$ with only the selected edges $\mathcal{E}_{T}$ together with the nodes $\mathcal{V}_{T}$ at both ends of each edge inside.
(2) Circuit fine-tuning: For a batch of data $\mathcal{X}_{T}\in\mathcal{X}$, all the parameters outside $\mathcal{C}$ are frozen, and only the parameters corresponding to the nodes $\mathcal{V}_{T}$ (see Section 3.2) are updated through a forward pass and a backward pass as usual. After $K$ steps of optimization, all the frozen parameters are freed and the graph $\mathcal{G}$ is reset to its original state.
In practice, it is efficient to accelerate graph pruning with only one forward and backward pass using edge attribution patching [56]. The implementation details are shown in Algorithm 1 and Section 5.
@@ 4.1 Circuit-tuning
  Input: dataset $\mathcal{X}$, model $M$, loss function $\mathcal{L}$, the metric for the quantification of the output $\mathcal{L}_{m}$, the number of edges to save $N$, the optimization steps after circuit discovery $K$.
  Initialize graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ from $M$, circuit $\mathcal{C}=\mathcal{G}$, and the iteration step $i=0$.
@@
           Intervene in edge $e$ and calculate the contribution of it according to Equation (7) and (D)
@@
        $\mathcal{C}\leftarrow\{\mathcal{V}_{T},\mathcal{E}_{T}\}$, where $\mathcal{V}_{T}=\!\{n\,|\,n\!\in\!\mathcal{V}\land n\text{ is incident to an edge }e\!\in\!\mathcal{E}_{T}\}$
@@
     Update the parameters corresponding to the nodes in $\mathcal{C}$
@@ 4.2 Characteristics of Our Method
Heuristic and dynamic circuit discovery Unlike prior works that focus on the circuits in a trained model, circuit-tuning acts as a heuristic approach for a model to develop an unseen ability in a dynamic style. The validity of this idea is built on the finding in [1] that pre-training has learned enough knowledge for downstream tasks, allowing fine-tuning with minor modifications.
Interpretable model diffing Instead of being a blackbox, circuit-tuning performs fine-tuning on a visible circuit, making the full process open to researchers. The transparency of our method provides the convenience to track the learning dynamics and the availability to intervene in the training process. Thus we are able to compare among different training stages and gain a deeper understanding and better control of it. Visualization of the training process is detailed in Section 5 and Appendix F.4.
Another feature of our method is that it is able to precisely locate the key components and update only a few parameters as discussed in Section 3.4, instead of modifying all of the parameters as LoRA-based methods do. Besides, while ReFT [64] a similar method also leverages representations for training, it is aimed at parameter-efficient fine-tuning and learns a transformation as an intervention on a hidden state, while our method is more focused on the mechanisms of learning during training.
5 Experiments
In this section, we test our theory and method across a variety of models and tasks. To comprehensively investigate the learning dynamics during fine-tuning as well as the practicality of our method, we arrange the experiments into two parts. The first part serves as the verification and analysis of our method, while the second part aims to show that our method is applicable to various scenarios.
5.1 The subject-verb disagreement Task
5.1.1 Task description and data preparation
The goal of the subject-verb disagreement task is to match a verb with a subject in an abnormal way. For example, “I is", “he are" and “the cows eats" are all expected results for this task. In each sentence, the token before the verb is called the END token. The automatic evaluation metric for this task is the logit difference between the flipped verb and the original verb at the END token. This is because the logit at the END token is directly used for predicting the verb.
The reason why we create this task as the start of our experiments is that we expect the model to acquire a new capability from scratch, and thus to study how the circuit evolves during fine-tuning. The task is simple and interesting for analyzing the learning dynamics with clear visualization.
Different from previous works [14, 34] that use template-based datasets, we collect real-world data from the Pile corpus [17] to ensure diversity and authenticity. We extract 60k sentences in the present tense in English and flip the forms of the verbs. We ensure the high quality of our dataset and believe it is meaningful for further research. For the details of our dataset, please refer to Appendix F.1.
5.1.2 Implementation
We use GPT2-small [47] in this task. We set the output of the attention and the MLP in each layer as upstream nodes, and the input of the query, key, value and the MLP in each layer as downstream nodes. During training, we use the logit difference discussed before as the metric $\mathcal{L}_{m}$. We follow Syed et al. [56] for path patching with mean ablation, with an improvement detailed in Appendix F.4.5. We sweep over a range of $N$ which is the number of edges to be saved during circuit discovery. Mini-batch SGD is used for optimization. We train the model under each setting for 3 epochs where the performance almost converges. For implementation details, please refer to Appendix F.2.
@@ 5.1.2 Implementation
Figure 1: Results in the subject-verb disagreement task.
@@ 5.1.3 Main results
From Figure 1(a), we observe a flip in logit difference from negative to positive, which means the model adjusts its grammar to fit the data distribution of subject-verb disagreement. One noteworthy finding is that the model can generate abnormal texts in the past tense, such as “to be or not to be, that were a …”, which implies that the model really grasps the new grammar and applies it smartly, since there is no sentence in the present tense in our training data at all.
@@
With the increasing number of top $N$ edges, the logit difference increases until $N$ approaches 1000. The number of tunable parameters is also saturated at this point. We believe that when $N=1000$, almost all the necessary parameters for this task are included, and the performance cannot be further improved, suggesting the minimality of the circuit and the existence of intrinsic dimensionality.
@@
We also track the perplexity (PPL) that is calculated over the test set with corpora in various scenarios. In Figure 1(b), we find that the PPL of full fine-tuning is high and fluctuates wildly during training, though the logit difference of it is higher. This implies that circuit-tuning provides better training stability as well as better preservation of the basic language modeling ability over full fine-tuning.
(a) Average tunable parameter ratio
(b) Final logit difference
(c) Final PPL
Figure 2: The changes in indexes with various choices of top-N imply the existence of the node-level intrinsic dimensionality at $N=1000$ in a computational graph. Note that the parameter ratio varies and depends on the granularity of nodes. Theoretically more granular nodes result in lower parameter ratio because the location of the intrinsic nodes would be more precise.
5.1.4 Ablation study for the quality of the discovered circuit
To confirm the effectiveness of our method to find the required parameters, we (i) calculate the faithfulness and completeness [62] of the circuits and (ii) provide an ablation study to randomly unfreeze the nodes outside the circuit during training and check if the performance is influenced. Results of the above experiments in Appendix F.3 ensure that the circuits are discovered accurately. Besides, we observe an obvious turning point of the faithfulness and completeness in Figure 8 as $N$ increases. Also, we notice a similar and common turning point of logit difference, PPL and number of tunable parameters in Figure 2. All of the above observations show the shadow of intrinsic nodes.
5.1.5 Analyses on learning dynamics
(a) Illustration of the flip.
(b) Before fine-tuning
(c) After fine-tuning
Figure 3: The flip observed in Finding 2.
Finding 1: Interpretation of the circuits
@@ 5.1.5 Analyses on learning dynamics
Inspired by Wang et al. [62], we decompose the circuit into attention heads of different functions. Firstly we find out the heads that directly affect the output, namely the Subject Attribute Heads. Then we find out the heads responsible for the localization of the subject, namely the Subject Identification Heads. Finally, we find out heads that could affect the behaviors of the Subject Attribute Heads, namely the Collaborative Heads. For technique details and visualization of the heads, please refer to Appendix F.4.1.
Finding 2: The flip of the Subject Attribute Heads
@@
The Subject Attribute Heads are responsible for matching the subject and the verb. We compute the dot product between the output $x$ of each attention head at the END token and $W_{U}(v_{flip})\!-\!W_{U}(v)$, the difference between the unembedding projections. Since the latter is fixed, we expect the projection of $x$ on the direction of it to be large, thus encouraging the probability difference between the two opposite verb forms (love v.s. loves, etc.). We visualize the heads before and after fine-tuning in Figure 3. Through comparison, we observe an obvious flip at head.10.9, implying the reversal in its function from subject-verb agreement to disagreement. Other heads (3.0, 6.0, 4.4, 11.8, etc) also see flips with varying degrees, serving as evidence for the self-adjustment of the functions inside the nodes. See Appendix F.4.1 for details.
Table 1: Some of the strengthened and weakened edges during training. The dynamic change shows the change in the logarithm of the edge contribution $log[1+c(e)]$ (the start of training $\rightarrow$ 2000 steps $\rightarrow$ 3000 steps $\rightarrow$ 4000 steps). The dynamic process is visualized in Figure 4.
Strengthened Edges Weakened Edges
Edge Dynamic change Edge Dynamic change
mlp.2 $\rightarrow$ mlp.3 0 $\rightarrow$ 0.0293 $\rightarrow$ 0.0993 $\rightarrow$ 0.2282 mlp.0 $\rightarrow$ mlp.10 0.2712 $\rightarrow$ 0.0580 $\rightarrow$ 0 $\rightarrow$ 0
(a) circuit at 2000 steps
(b) circuit at 3000 steps
(c) circuit at 4000 steps
Figure 4: Evidence of Hebbian learning in the subject-verb disagreement task ($N=50$ edges shown).
Finding 3: The sharing of the Subject Identification Heads
@@
The Subject Identification Heads attend heavily to the subject. One type of these heads attends to the END token (0.1, 0.3, etc), which is helpful to the cases like “they are"; the other type of heads (8.5, 10.5, etc) attends to the subject several tokens before, which is helpful to the cases like “the girl wearing a dress is". Both types of heads perform the same before and after fine-tuning. This implies that their functions are preserved and shared all the way down. Our finding is consistent with the study of intrinsic dimensionality in Aghajanyan et al. [1] that pre-training optimizes the description length without having direct access to the same tasks. The kind of consistency further confirms the feasibility of our theory in Section 3.
Finding 4: The interaction and collaboration inside the model
Interestingly, we observe that over the training process the nodes (attention heads) inside a circuit complete a task through a cooperative division of labor. Each head has its division of labor as discussed before. Besides, some heads directly affect the output, while others cooperate with them to affect the output indirectly. Several heads achieve a common goal through cooperation: they could affect a head only through combined effect, and adjust themselves during training. For details, please refer to Appendix F.4.1.
Finding 5: The evidence of Hebbian learning
@@
We visualize the circuits during the fine-tuning process in Figure 4, and observe a phenomenon akin to Hebbian learning [10] that several edges are strengthened during training, just as the synapses between neurons can be strengthened after continuous stimulation. Also, several edges are weakened at the same time, as shown in Table 1. It is similar to the lateral inhibition among neurons [22] that an activated neuron can reduce the activity of its neighbors. We present the dynamics of the competitive learning and self-organization [25] inside a model, providing new discoveries to the generalization process on a new data distribution.
@@
Due to limited space, all the above together with other interesting findings are detailed in F.4.
5.2 Application to Complex Tasks